I need help on how to set up the word problems. The sum of three is one-hundred twenty-three. the second number is two more than twice the first number . the third is five less than the produce of three and the first number. find the three numbers. I understand the total is 123 I just don't understand the rest can you please help me with the steps if you can't can you tell me a site that I can go to to help me understand. Thank Kary Martin : I need help on how to set up the word problems. The sum of three is one-hundred twenty-three. the second number is two more than twice the first number . the third is five less than the produce of three and the first number. find the three numbers. I understand the total is 123 I just don't understand the rest can you please help me with the steps if you can't can you tell me a site that I can go to to help me understand. Thank Kary Martin
Answer by ankor@dixie-net.com(4532)

Let the 3 numbers be: x, y z
:
Write and equation for each statement:
:
"The sum of three numbers is one-hundred twenty-three."
x + y + z = 123
:
" the second number is two more than twice the first number:
y = 2x + 2
:
"the third is five less than the product of three and the first number."
z = 3x - 5
;
find the three numbers.
:
Substitute for y and z in the total equation
:
x + (2x+2) + (3x-5) = 123
:
6x - 3 = 123
:
6x = 123 + 3
:
6x = 126
x = 126/6
x = 21
:
Use the other two equations to find y and z, check the total
:
has this been set up enough for you??
Question 153147: I need help on how to set up the word problems. The sum of three is one-hundred twenty-three. the second number is two more than twice the first number . the third is five less than the produce of three and the first number. find the three numbers. I understand the total is 123 I just don't understand the rest can you please help me with the steps if you can't can you tell me a site that I can go to to help me understand. Thank Kary Martin : I need help on how to set up the word problems. The sum of three is one-hundred twenty-three. the second number is two more than twice the first number . the third is five less than the produce of three and the first number. find the three numbers. I understand the total is 123 I just don't understand the rest can you please help me with the steps if you can't can you tell me a site that I can go to to help me understand. Thank Kary Martin
Answer by stanbon(18999)

The sum of three is one-hundred twenty-three. the second number is two more than twice the first number . the third is five less than the produce of three and the first number. find the three numbers.
----------------------
EQUATIONS:
a + b + c = 123
b = 2a +2
c = 3a - 5
-----------------
Rearrange to get:
a + b + c = 123
2a - b + 0 = -2
3a + 0 - c = 5
------------------------
Use any method to solve the system to get:
a = 21
b = 44
c = 58
============
Cheers,
Stan H.